# Using Knowledge Graphs to harvest datasets for efficient CLIP model training

### v1

@@ 4.4 Retrieval evaluation
We evaluate the COCO Karpathy test split [15], a subset of 5000 samples from the MS-COCO [19] dataset paired with 5 text each. We also evaluate the 1000 samples in the Karpathy test split of Flickr30k [50] annotated with 5 texts per image, as well as on the 3600 image-text pairs in XM3600 [37]. We report the average of image-to-text and text-to-image recall@1 over all datasets.
@@ 5 Results
Figure 3: Results on image retrieval and robustness against distribution shifts on ImageNet.
Table 4: Results for training CLIP B-32 and B-16 on our EntityNet dataset from scratch. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training. We only compare zero-shot results and mark results as “–” if the model does not fulfill the zero-shot requirements.
@@
Table 5: Results for finetuning the DataComp-1B CLIP model on the EntityNet dataset. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training.
@@
B-32 EntityNet (only living organisms) 4.2 10.2 48.2 31.2 33.3 37.0 87.0
B-16 DataComp-1B (Base model) 791.4 1400.0 73.5 57.4 64.4 15.3 79.0
@@
B-16 EntityNet (only living organisms) 11.3 10.2 51.4 34.8 39.2 42.7 90.3
Table 6: Analysis of performance when varying dataset composition, text sampling and dataset size. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training.
@@
B-32 No living organisms 9.0 22.5 39.2 32.1 28.0 0.8 6.2 6.9
B-32 Only living organisms 4.1 10.2 36.0 16.5 21.0 28.6 83.2 42.0
B-32 No attribute queries 8.7 21.8 54.8 28.6 33.8 25.6 79.7 39.2
@@
We evaluate CLIP pretrained from scratch on our EntityNet dataset and CLIP models trained on other datasets. In Figures 1 and 3 we contrast the effectiveness of models with their training cost. We show the results in detail in Table 4. In the generic domain, our models surpass others trained on similarly sized datasets while achieving comparable performance on object classification with models trained 20x longer. On image-text retrieval, our model performs similarly to models trained on the same amount of compute. While our pipeline creates a dataset efficient for understanding objects and their properties, understanding complex scenes still requires learning mainly from the alt texts more than from objects and attributes. In the expert domain, we outperform even the largest CLIP models on the challenging iNaturalist 2021 dataset, which requires classifying images among 10k fine-grained species. Our model also excels on CUB by distinguishing 200 bird species better than all other CLIP models of the same size. Further, when compared to the expert model BioCLIP, trained specifically for organismal biology at a similar training cost, our model demonstrates superior performance. On the Rare Species benchmark, our model classifies the species unseen during training better than the expert biology model, showing the effectiveness of our dataset collection method over a manually designed living organism training set.
We investigate improving existing CLIP models via finetuning in Table 5. The results show that our dataset can be leveraged to create expert CLIP models that outperform both the base model and our model pretrained from scratch on the expert domain. This improvement comes at the cost of trading off some capabilities in the other domains. When finetuning only on the expert domain, we trade off more capabilities, yet obtain even stronger experts.
We further validate and verify our design choices through a component analysis in Table 6. Training separately on the generic and the domain expert part of our dataset reveals that, while the best generic model emerges from training on everything, a slightly better expert model is the result of training only on the expert domain (first table segment). However, generalization to unseen species slightly benefits when training on the full dataset, showing that our generic domain data can enhance generalization capabilities within the expert domain. We also observe that generating and downloading attribute queries contributes to improved performance of the pretrained model.
@@ 6 Limitations
The proposed data harvesting approach assumes that there is a knowledge graph for the target domain and that there is a searchable database with noisy pairing of images and text. Especially, the latter assumption can induce extra effort in some domains. In the medical domain, for example, there are massive amounts of paired image-text data, but the data is not publicly available and not connected to a regular search engine. The search queries must be adapted appropriately to the respective local database. Another limitation is the small, but significant drop in performance on image-text retrieval and classifying ImageNet distribution shifts in the generic domain, when finetuning a large model with EntityNet. First, our dataset by design has a strong focus on the expert domain and trades off some performance in the generic domain during finetuning. Second, our search pipeline finds many clean object-centric images and annotates them with entity information, which tremendously helps understanding object semantics, but to improve efficiency on image-text retrieval in a similar way, one needs to tackle the quality of alt texts and their alignment to the images [47]. Finally, we focused on searching photos, which explains slightly lower accuracy when classifying paintings and sketches – the EntityNet dataset simply contains a lower percentage of such types of images than datasets like CommonPool.
@@ 7 Conclusions
We demonstrated how to use knowledge graphs to harvest datasets that are efficient for training CLIP models. Our strategy allows us to create an expert domain dataset with little manual effort, enabling the development of CLIP models that significantly outperform standard models in the expert domain. The expert domain dataset can be used either for training a model from scratch or for finetuning an existing vanilla model. The substantial size and diversity of the expert domain dataset ensures that the good generalization properties of CLIP exist also in the expert domain, in contrast to training with an over-specialized expert dataset.
@@
A Detailed results
@@
We show extensive results on object classification in Table 7. For a more detailed analysis of model capabilities on ImageNet, we split the classes into living (410 classes) and other (590 classes) using WordNet: Since ImageNet labels are built on WordNet nouns, we simply select all labels that are children of the living things node for the living set. On iNaturalist, in addition to the 2021 version, we also evaluate on the 2019 version which contains 3,030 images in the validation set, each annotated with one of 1,010 species. We test with the same protocol as on iNaturalist 2021, testing on both english and latin class names and reporting the best accuracy. We show additional results on retrieval in Table 8 and object classification under distribution shift in Table 9.
@@
Table 7: Detailed object classification results. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training. We only compare zero-shot results and mark results as “–” if the model does not fulfill the zero-shot requirements.
Arch. Dataset MACs Images in ImageNet iNaturalist CUB Rare
(1e18) dataset 1k Living other 2019 2021 species
[…]
@@
B-32 $\star$ EntityNet (only living organisms) 4.2 10.2 48.2 76.3 30.9 49.3 37.0 87.0
B-16 DataComp-1B (Base model) 791.4 1400.0 73.5 75.9 73.2 19.5 15.3 79.0
@@
B-16 $\star$ EntityNet (only living organisms) 11.3 10.2 51.4 80.2 33.8 54.3 42.7 90.3
B-32 EntityNet Everything 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 EntityNet (no living organisms) 9.0 22.5 39.2 17.5 56.1 1.7 0.8 6.2 6.9
B-32 EntityNet (only living organisms) 4.1 10.2 36.0 68.5 15.4 41.4 28.6 83.2 42.0
B-32 EntityNet (no attribute queries) 8.7 21.8 54.8 63.0 50.4 36.4 25.6 79.7 39.2
B-32 EntityNet 50% alt text 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 EntityNet 100% alt text 13.1 32.7 59.1 66.5 55.4 36.4 22.9 78.8 39.7
B-32 EntityNet 0% alt text 13.1 32.7 55.7 64.4 51.5 35.4 24.2 78.1 29.7
B-32 EntityNet Full size 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 EntityNet 1/2 size 6.6 16.4 54.1 61.5 50.5 30.7 20.1 74.1 36.6
B-32 EntityNet 1/4 size 3.3 8.2 45.2 52.4 41.8 23.4 13.2 64.9 28.0
B-32 EntityNet 1/8 size 1.6 4.1 33.3 39.6 30.5 14.5 7.3 47.8 19.4
B-32 EntityNet 1/16 size 0.8 2.0 19.9 25.0 18.1 7.6 2.8 27.4 11.8
B-32 EntityNet batch size 2,048 13.1 32.7 59.7 65.4 57.2 31.5 21.0 74.1 38.7
B-32 EntityNet batch size 4,096 13.1 32.7 60.9 67.0 58.0 36.2 24.1 77.6 41.0
B-32 EntityNet batch size 8,192 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 EntityNet batch size 16,384 13.2 32.7 60.5 68.2 56.6 38.8 26.9 81.2 42.1
B-32 EntityNet batch size 32,768 13.3 32.7 58.6 67.0 54.2 39.5 26.3 81.3 41.7
Table 8: Detailed retrieval results. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training.
Arch. Dataset MACs Images in Retrieval COCO test F30K test XM3600 test
(1e18) dataset Average I2T T2I I2T T2I I2T T2I
@@
B-32 $\star$ EntityNet (only living organisms) 4.2 10.2 31.2 28.0 19.7 44.8 33.7 30.9 29.9
B-16 DataComp-1B (Base model) 791.4 1400.0 57.4 57.5 40.2 84.9 67.3 47.9 46.5
@@
B-16 $\star$ EntityNet (only living organisms) 11.3 10.2 34.8 31.8 22.8 51.7 37.1 33.5 32.0
B-32 EntityNet Everything 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 EntityNet (no living organisms) 9.0 22.5 32.1 28.6 18.7 46.3 33.6 33.2 32.3
B-32 EntityNet (only living organisms) 4.1 10.2 16.5 12.8 9.8 23.4 15.3 19.0 18.6
B-32 EntityNet (no attribute queries) 8.7 21.8 28.6 23.6 16.1 41.3 27.3 32.3 30.8
B-32 EntityNet 50% alt text 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 EntityNet 100% alt text 13.1 32.7 38.3 35.2 23.2 53.4 39.8 40.5 38.0
B-32 EntityNet 0% alt text 13.1 32.7 13.5 8.7 6.1 19.2 11.7 18.3 17.0
B-32 EntityNet Full size 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 EntityNet 1/2 size 6.6 16.4 30.3 27.0 17.6 42.7 29.7 33.3 31.6
B-32 EntityNet 1/4 size 3.3 8.2 23.5 19.5 13.0 31.1 22.9 27.6 27.2
B-32 EntityNet 1/8 size 1.6 4.1 16.6 12.8 8.7 20.7 14.2 21.9 20.9
B-32 EntityNet 1/16 size 0.8 2.0 9.4 7.1 5.0 10.0 7.1 13.9 13.5
B-32 EntityNet batch size 2,048 13.1 32.7 35.6 30.8 20.9 50.2 37.2 38.2 36.4
B-32 EntityNet batch size 4,096 13.1 32.7 36.4 31.6 22.1 51.7 37.0 39.1 37.2
B-32 EntityNet batch size 8,192 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 EntityNet batch size 16,384 13.2 32.7 35.8 32.2 21.8 50.5 36.9 37.2 36.1
B-32 EntityNet batch size 32,768 13.3 32.7 34.5 31.0 20.8 48.9 34.5 36.2 35.3
Table 9: Detailed results on ImageNet distribution shifts. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training. INet: ImageNet.
Arch. Dataset MACs Images in INet Ave- INet INet Object- INet INet
@@
B-32 $\star$ EntityNet (only living organisms) 4.2 10.2 48.2 33.3 43.1 56.7 19.1 32.1 15.8
B-16 DataComp-1B (Base model) 791.4 1400.0 73.5 64.4 66.0 83.6 63.7 60.4 48.4

### v3

@@ 4.4 Retrieval evaluation
We evaluate the COCO Karpathy test split [46], a subset of 5000 samples from the MS-COCO [47] dataset paired with 5 texts each. We also evaluate the 1000 samples in the Karpathy test split of Flickr30k [48] annotated with 5 texts per image, as well as on the 3600 image-text pairs in XM3600 [49]. We report the average of image-to-text and text-to-image recall@1 over all datasets.
@@ 5 Results
Figure 3: Results on image retrieval and distribution shift robustness on ImageNet.
Table 3: Results for training CLIP B-32 and B-16 on our EntityNet dataset from scratch. We mark the best and second best result. To measure compute cost, we calculate training MACs (multiply–accumulate operations). We only compare zero-shot results and mark results as “–” if the model does not fulfill the zero-shot requirements.
@@
Table 4: Results for finetuning the DataComp-1B CLIP model on EntityNet. We mark the best and second best result. To measure compute cost, we calculate training MACs (multiply–accumulate operations).
@@
B-32 Only organisms 4.2 10.2 48.2 31.2 33.3 37.0 87.0
B-16 DataComp-1B 791.4 1400.0 73.5 57.4 64.4 15.3 79.0
@@
B-16 Only organisms 11.3 10.2 51.4 34.8 39.2 42.7 90.3
Table 5: Analysis of performance when varying dataset composition, text sampling and dataset size. We mark the best and second best result. To measure compute cost, we calculate training MACs (multiply–accumulate operations).
@@
B-32 No organisms 9.0 22.5 39.2 32.1 28.0 0.8 6.2 6.9
B-32 Only organisms 4.1 10.2 36.0 16.5 21.0 28.6 83.2 42.0
B-32 No attributes 8.7 21.8 54.8 28.6 33.8 25.6 79.7 39.2
@@
We evaluate CLIP pretrained from scratch on our EntityNet dataset and CLIP models trained on other datasets. In Figures 1 and 3 we contrast the effectiveness of models with their training cost. We show the results in detail in Table˜3. In the generic domain, our models surpass others trained on similarly sized datasets while achieving comparable performance on object classification with models trained 20x longer. On image-text retrieval, our model performs similarly to models trained on the same amount of compute. While our pipeline creates a dataset efficient for understanding objects and their properties, understanding complex scenes still requires learning mainly from the alt texts more than from objects and attributes. In the expert domain, we outperform even the largest CLIP models on the challenging iNaturalist 2021 dataset, which requires classifying images among 10k fine-grained species. Our model also excels on CUB by distinguishing 200 bird species better than all other CLIP models of the same size. Further, when compared to the expert model BioCLIP, explicitly trained for organismal biology at a similar training cost, our model demonstrates superior performance. On the Rare Species benchmark, our model outperforms BioCLIP on unseen species, showing the effectiveness of our dataset collection method over a manually designed living organism training set.
We investigate improving existing CLIP models via finetuning in Table˜4. The results show that our dataset can be leveraged to create expert CLIP models that outperform both the base model and our model pretrained from scratch on the expert domain. This improvement comes at the cost of trading off some capabilities in the other domains. When finetuning only on the expert domain, we trade off more capabilities, yet obtain even stronger experts.
We further validate and verify our design choices through a component analysis in Table˜5. Training separately on the generic and the domain expert part of our dataset reveals that, while the best generic model emerges from training on everything, a slightly better expert model is the result of training only on the expert domain (first table segment). However, generalization to unseen species slightly benefits when training on the full dataset, showing that our generic domain data can enhance generalization capabilities within the expert domain. We also observe that generating and downloading attribute queries contributes to improved performance of the pretrained model.
@@ 6 Limitations
The proposed data harvesting approach assumes that there is a knowledge graph for the target domain and that there is a searchable database with noisy pairing of images and text. However, knowledge graphs exist in many domains, e.g., UniProt [50] with 246M protein sequence records or AgriKG [51] with 150k agricultural entities. Also, if no image search engine is available for the given domain, but a large amount of image-text data exists, pairs can be found by searching for the queries via substring matching in the image-text pairs.
Another limitation is the small, but significant drop in performance on image-text retrieval and classifying ImageNet distribution shifts in the generic domain, when finetuning a large model with EntityNet. First, our dataset by design has a strong focus on the expert domain and trades off some performance in the generic domain during finetuning. Second, our search pipeline finds many clean object-centric images and annotates them with entity information, which tremendously helps understanding object semantics, but to improve efficiency on image-text retrieval in a similar way, one needs to tackle the quality of alt texts and their alignment to the images [52]. Finally, we focused on searching photos, which explains slightly lower accuracy when classifying paintings and sketches – the EntityNet dataset simply contains a lower percentage of such types of images than datasets like CommonPool.
@@ 7 Conclusions
We demonstrated how to use knowledge graphs to harvest datasets that are efficient for training CLIP models. Our strategy allows to create an expert domain dataset with little manual effort, enabling the development of CLIP models that significantly outperform standard models in the expert domain. The expert domain dataset can be used for training a model from scratch or for finetuning an existing vanilla model. The substantial size and diversity of the expert domain dataset ensures that the good generalization properties of CLIP exist also in the expert domain, in contrast to training with an over-specialized expert dataset.
@@
Acknowledgements
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499552394 – SFB 1597 – Project-ID 417962828 – Project-ID 539134284. The authors acknowledge support from the state of Baden-Württemberg through bwHPC.
A Additional analyses
A.1 Verification of Bing search results
In this work, we search for images that fit our queries using Bing image search. In contrast, CLIP [1] searches for the query in a large pool of image-text pair candidates created from raw HTML via substring matching. We manually evaluate the quality of the image search engine and of substring matching in Table˜6, using randomly select queries of our EntityNet dataset. In total, 87% of queries were answered correctly. Incorrect results of Bing search can be grouped into wrong similar terms (for the japanese “dogi” uniform, the search instead returns actual dogs), and attribute ignored (for “B-25 Mitchell gray paint” the search returns the correct B-25 plane, but differently painted). Bing search is clearly superior to substring matching, with the latter algorithm only fulfilling 52% of queries.
A.2 Scaling behaviour of EntityNet
In Figure˜4, We study the impact of removing entire entities or reducing the number of queries per entity, and compare this with randomly dropping images. Removing entities and their associated images degrades performance more than randomly removing the same number of images, indicating that entities represent important semantic concepts for model training. In other words, our results suggest that diversity (more entities with fewer images per entity) is more important than depth (fewer entities with more images each). In all cases, shrinking the dataset lowers performance, so a certain dataset size (here 33M) is indeed essential for a good generic model.
B Detailed experimental results
@@
We show extensive results on object classification in Table˜7. For a more detailed analysis of model capabilities on ImageNet, we split the classes into living (410 classes) and other (590 classes) using WordNet: Since ImageNet labels are built on WordNet nouns, we simply select all labels that are children of the living things node for the living set. On iNaturalist, in addition to the 2021 version, we also evaluate on the 2019 version which contains 3,030 images in the validation set, each annotated with one of 1,010 species. We test with the same protocol as on iNaturalist 2021, testing on both english and latin class names and reporting the best accuracy. We show additional results on retrieval in Table˜8 and object classification under distribution shift in Table˜9.
C Experimental details
C.1 Hyperparameter settings
We report hyperparameters for our experiments in Table˜10. Similar to Li et al. [23], we reduce the context size of the text encoder down from 77 to 32 to reduce VRAM and training time requirements. For a fair comparison with other CLIP models, we report all training cost and training duration as if the training was run at a context length of 77.
C.2 Object classification evaluation
To test the VLMs on object classification, we use the same procedure as CLIP [1]: Given an image $I$, class names $C_{1},...,C_{N}$, image encoder $f$ and text encoder $g$, we embed the image using the image encoder $\mathbf{v}=f(I)$. To acquire a text embedding for class $C_{c}$, the CLIP authors started by directly encoding the class names as $\mathbf{w}_{c}=g(C_{c})$, e.g., dog. Alternatively, they created several prompts $P$ using templates, e.g., graffiti of a dog, a photo of the cool dog, etc., then encoded each prompt, and computed the average embedding: $\mathbf{w}_{c}=\sum_{p\in P}g(p)/|P|$. They referred to this approach as using “context prompts”. Finally, given the image and text embeddings, the prediction $p$ is the class which has the highest cosine similarity to the image.
For a fair comparison between models that have been trained with different prompts, we evaluate all models on just encoding the class name, and on using the average embedding of the 80 context prompts that the CLIP authors used for ImageNet, then report the higher top-1 accuracy.
D Qualitative example of text label sampling
We show an example of our text label sampling strategy in Table˜11.
E Image Search API details
Google The Google Image Search API is available via the Google Cloud Platform, and requires an existing programmable search engine to function. It returns up to 10 images per request and page with a limit of 10 pages, i.e., 100 images per query. It costs 5\$ per 1,000 API calls, resulting in costs of about 500\$ to download 1M images. We found the search results from the Google API to be quite different, and arguably worse, from the ones returned when using the regular Google image search. For all our API requests we set the parameter imgColorType to color, imgType to photo, lr to lang_en, and excludeTerms to drawing clipart illustration cartoon vector painting. This way we get mostly real-world images in our search results. We additionally add all aliases and the natural type of the sought entity to the orTerms parameter for entity and entity-attribute queries. Because the Google API returns only up to 10 images per request and page, we search for the following number of pages: 2 pages each for entity queries, 4 pages each for entity-attribute queries, and 10 pages each for all natural-type-attribute queries. We started our search with queries from the organism subset on both the Google and Bing APIs. We found the quality and value-for-money ratio of the Bing API to be better, and therefore switched to only using Bing for all other queries.
F Querying entities with SPARQL
The SPARQL query over Wikidata used to harvest all entities under a super-entity is displayed in Figure˜5. It returns a list of entities from a specified target domain as defined by one or more super-entities. The super-entities can be determined manually by searching for appropriate entities on the Wikidata website. For example, if we want to build a dataset about vehicles, we can set the super-entity to vehicle (Q42889), as done in Table˜12. We list the super-entities we considered for our dataset and the relevant statistics in Table˜13 for included and Table˜14 for excluded super-entities.
Table 6: We randomly select 100 search queries for each query set and manually check for the following errors: Wrong: The majority of the results do not match the query. Too few: Only four or less images are found. Over all 416k queries, Bing search finds $\geq 5$ images in 99.8% of cases.
Query set Bing search Substring matching
Wrong Too few Wrong Too few
World entity 13% 0% 15% 26%
World entity + attribute 22% 0% 3% 66%
Living entity 0% 0% 7% 19%
Living entity + attribute 18% 0% 7% 51 %
Total 13% 0% 8% 40 %
Figure 4: Scaling entities and queries per entity.
@@ F Querying entities with SPARQL
Table 7: Detailed object classification results. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure compute cost, we calculate training MACs (multiply–accumulate operations). We only compare zero-shot results and mark results as “–” if the model does not fulfill the zero-shot requirements.
Arch. Dataset MACs Images in ImageNet iNat. CUB Rare
(1e18) dataset (M) 1k living other 2019 2021 spcs
# Classes $\rightarrow{}$ 1,000 410 590 1,010 10k 200 400
@@
B-32 $\star$ Only organisms 4.2 10.2 48.2 76.3 30.9 49.3 37.0 87.0
B-16 DataComp-1B 791.4 1400.0 73.5 75.9 73.2 19.5 15.3 79.0
@@
B-16 $\star$ Only organisms 11.3 10.2 51.4 80.2 33.8 54.3 42.7 90.3
B-32 EntityNet 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 No organisms 9.0 22.5 39.2 17.5 56.1 1.7 0.8 6.2 6.9
B-32 Only organisms 4.1 10.2 36.0 68.5 15.4 41.4 28.6 83.2 42.0
B-32 No attributes 8.7 21.8 54.8 63.0 50.4 36.4 25.6 79.7 39.2
B-32 50% alt text 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 100% alt text 13.1 32.7 59.1 66.5 55.4 36.4 22.9 78.8 39.7
B-32 0% alt text 13.1 32.7 55.7 64.4 51.5 35.4 24.2 78.1 29.7
B-32 Full size 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 1/2 size 6.6 16.4 54.1 61.5 50.5 30.7 20.1 74.1 36.6
B-32 1/4 size 3.3 8.2 45.2 52.4 41.8 23.4 13.2 64.9 28.0
B-32 1/8 size 1.6 4.1 33.3 39.6 30.5 14.5 7.3 47.8 19.4
B-32 1/16 size 0.8 2.0 19.9 25.0 18.1 7.6 2.8 27.4 11.8
B-32 Batch size 2k 13.1 32.7 59.7 65.4 57.2 31.5 21.0 74.1 38.7
B-32 Batch size 4k 13.1 32.7 60.9 67.0 58.0 36.2 24.1 77.6 41.0
B-32 Batch size 8k 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 Batch size 16k 13.2 32.7 60.5 68.2 56.6 38.8 26.9 81.2 42.1
B-32 Batch size 32k 13.3 32.7 58.6 67.0 54.2 39.5 26.3 81.3 41.7
Table 8: Detailed retrieval results. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure compute cost, we calculate training MACs (multiply–accumulate operations).
Arch. Dataset MACs Images in Ave- COCO Flickr30K XM3600
(1e18) dataset (M) rage I2T T2I I2T T2I I2T T2I
@@
B-32 $\star$ Only organisms 4.2 10.2 31.2 28.0 19.7 44.8 33.7 30.9 29.9
B-16 DataComp-1B 791.4 1400.0 57.4 57.5 40.2 84.9 67.3 47.9 46.5
@@
B-16 $\star$ Only organisms 11.3 10.2 34.8 31.8 22.8 51.7 37.1 33.5 32.0
B-32 EntityNet 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 No organisms 9.0 22.5 32.1 28.6 18.7 46.3 33.6 33.2 32.3
B-32 Only organisms 4.1 10.2 16.5 12.8 9.8 23.4 15.3 19.0 18.6
B-32 No attributes 8.7 21.8 28.6 23.6 16.1 41.3 27.3 32.3 30.8
B-32 50% alt text 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 100% alt text 13.1 32.7 38.3 35.2 23.2 53.4 39.8 40.5 38.0
B-32 0% alt text 13.1 32.7 13.5 8.7 6.1 19.2 11.7 18.3 17.0
B-32 Full size 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 1/2 size 6.6 16.4 30.3 27.0 17.6 42.7 29.7 33.3 31.6
B-32 1/4 size 3.3 8.2 23.5 19.5 13.0 31.1 22.9 27.6 27.2
B-32 1/8 size 1.6 4.1 16.6 12.8 8.7 20.7 14.2 21.9 20.9
B-32 1/16 size 0.8 2.0 9.4 7.1 5.0 10.0 7.1 13.9 13.5
B-32 Batch size 2k 13.1 32.7 35.6 30.8 20.9 50.2 37.2 38.2 36.4
B-32 Batch size 4k 13.1 32.7 36.4 31.6 22.1 51.7 37.0 39.1 37.2
B-32 Batch size 8k 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 Batch size 16k 13.2 32.7 35.8 32.2 21.8 50.5 36.9 37.2 36.1
B-32 Batch size 32k 13.3 32.7 34.5 31.0 20.8 48.9 34.5 36.2 35.3
Table 9: Detailed results on ImageNet distribution shifts. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure compute cost, we calculate training MACs (multiply–accumulate operations). INet: ImageNet.
Arch. Dataset MACs Images in INet Ave- INet INet Obj.- INet INet
@@
B-32 $\star$ Only organisms 4.2 10.2 48.2 33.3 43.1 56.7 19.1 32.1 15.8
B-16 DataComp-1B 791.4 1400.0 73.5 64.4 66.0 83.6 63.7 60.4 48.4